OpenStudy (cbreadbox):

I have no idea how to solve this question. "Find the work done by a 10 pound force acting in the direction (1, 2) in moving an object 3 feet from (0, 0) to (3, 0). "

OpenStudy (greatlife44):

\[F*dCos(\phi) = Work\]

OpenStudy (greatlife44):

where phi is the angle between the applied force and direction the object is moving.

OpenStudy (greatlife44):

I'm assuming that the applied force is being applied horizontally. if it were at an angle we would need to only consider the x component of the force since the object moves horizontally. well that's a different case I guess.  \[3~feet*(\frac{ 0.38m}{ft }) = 1.14~meters \] \[\phi = 0~degrees \] our applied force is 10 pounds, that needs to be in kg, and also force is in Newtons \[N = kg~\frac{ m }{ s^{2} }\] 10 pounds = 4.5kg \[4.5kg*9.8\frac{ m }{ s^{2} }*(1.14m)*Cos(0^{o}) = +50.3J\]